OpenStudy (anonymous):

After how many months will Rita pay two hundred dollars in simple annual interest on a credit card that charges fifteen percent on her four hundred dollars balance? 2 months fifteen months twenty-eight months forty months

OpenStudy (anonymous):

As given: P = $400 R=15%(whopping high rate man) \[T = \frac{ I \times 100 }{ P \times R} = 200 \times 100 / 400 \times 15 = \frac{ 50 }{ 15} = 3 months\] ..Not sure ..what's the correct answer?

OpenStudy (anonymous):

Maybe it can be 2 months?

OpenStudy (anonymous):

yes 2 months

OpenStudy (anonymous):

because when i did the math i got 2.3

OpenStudy (anonymous):

months

OpenStudy (anonymous):

Thanks Cram.

OpenStudy (anonymous):

welcome